Please help me solve this word problem on age: Six years ago, Wendy was twice the age of Thao. At present, Wendy is 30 years older than Thao. Find the present age of each woman.
Answer by mananth(16946)

Thao = x years now
Wendy = x+30
6 years ago
Thao = x-6
Wendy = (x+30)-6==>x+24
x+24=2(x-6)
x+24=2x-12
2x-x=24+12
x=36 years Thao's age now
36+30 = 66 years Wendy's age
m.ananth@hotmail.ca